Two thin concentric spherical shells of radii ri a Solve a, b and c
Physics
Answer #1

two concentric spherical conducting shells of radii r1,r2 have charges q1 and q2 respectively

case (1) if r<r1 (for points inside the inner shell)

Qenclosed =0

the electric field E*4\pi r^2=Qenclosed/\epsilon

E*4\pi r^2=0/\epsilon

the electric field E=0

case 2 if r1<r<r2( for points between the shells )

Qenclosed =q1

E*4\pi r^2=Qenclosed/\epsilon

E*4\pi r^2=q1/\epsilon

E*4\pi r^2=\sigma 1*4\pi r1^2/\epsilon

the electric field E=\sigma 1* r1^2/r^2\epsilon

case 3:if r>r2 (foe points outside the shell)

Qenclosed =q1+q2

the electric field E*4\pi r^2=Qenclosed/\epsilon

E*4\pi r^2=q1+q2/\epsilon

E*4\pi r^2=[\sigma 1*4\pi r1^2+\sigma 2*4\pi r2^2]/\epsilon

The electric field E=[\sigma 1* r1^2+\sigma 2* r2^2]/r^2\epsilon